# Taming Hyperparameter Sensitivity in Data Attribution: Practical Selection Without Costly Retraining

### v1

@@ 4.1 Analysis: Hyperparameter Selection of $\lambda$ in Influence Function
Remark 4.4 (Notation Simplification).
In Theorem 4.3, we have simplified the notation through the following definitions. We let $D_{a}$ denote the uniform distribution of size $a$ subsets of $S$. For $k\in\mathbb{N}^{*}$, denote $t_{k,z^{\prime},\lambda}:=\nabla_{\theta}f(z^{\prime},\theta_{S}^{*})^{\top}(F_{S}+\lambda I_{p})^{-k}F_{S}\nabla_{\theta}f(z^{\prime},\theta_{S}^{*})$. Define $g_{z^{\prime}}:=\frac{1}{n}\sum_{i=1}^{n}\alpha_{z^{\prime},i}\nabla_{\theta}L(z_{i},\theta_{S}^{*})\in\mathbb{R}^{p}$, where $\alpha_{z^{\prime}}\in\mathbb{R}^{n}$ with $\alpha_{z^{\prime},i}:=\mathbb{E}_{A\sim D_{a}}[f(z^{\prime},\theta_{A}^{*})|z_{i}\in A]-\mathbb{E}_{A\sim D_{a}}[f(z^{\prime},\theta_{A}^{*})]$. Further, we let $r_{z^{\prime},\lambda}:=-\nabla_{\theta}f(z^{\prime},\theta_{S}^{*})^{\top}(F_{S}+\lambda I_{p})^{-1}g_{z^{\prime}}$ and $o_{z^{\prime},\lambda}:=\alpha_{z^{\prime}}^{\top}(JJ^{\top}+n\lambda I_{n})^{-1}\alpha_{z^{\prime}}$, where $J\in\mathbb{R}^{n\times p}$ has its $i$th row being $\nabla_{\theta}L(z_{i},\theta_{S}^{*})^{\top}$.
@@
A result similar to Theorem 4.3 can also be derived for IFFIM attributor with gradient projection (see $\tau_{\mathrm{IFFIM},\lambda,P}$ in Eq. 4). In this case, given the projection matrix $P$, $t_{k,z^{\prime},\lambda}$ should be replaced with $\nabla_{\theta}f(z^{\prime},\theta_{S}^{*})^{\top}P(P^{\top}F_{S}P+\lambda I_{\tilde{p}})^{-k}P^{\top}F_{S}PP^{\top}\nabla_{\theta}f(z^{\prime},\theta_{S}^{*})$ while $r_{z^{\prime},\lambda}$ should be replaced with $-\nabla_{\theta}f(z^{\prime},\theta_{S}^{*})^{\top}P(P^{\top}F_{S}P+\lambda I_{\tilde{p}})^{-1}P^{\top}g_{z^{\prime}}$. Additionally, $JJ^{\top}+n\lambda I_{n}$ in the $o_{z^{\prime},\lambda}$ should be replaced with $JPP^{\top}J^{\top}+n\lambda I_{n}$. Please see more details in Section B.4.
@@ 4.2 Algorithm: The Surrogate-Indicator-Based Practical Selection
The sufficient condition in Theorem 4.3 enables a practical selection algorithm for $\lambda$ based on the following surrogate indicator $\xi_{z^{\prime},\lambda}$.
@@
$\xi_{z^{\prime},\lambda}:=\frac{t_{2,z^{\prime},\lambda}}{\sqrt{t_{3,z^{\prime},\lambda}\cdot t_{1,z^{\prime},\lambda}}},$ (9)
which is the RHS of Eq. 8.
Note that the quantities $t_{k,z^{\prime},\lambda}$ for $k=1,2,3$ only depends on model parameters $\theta_{S}^{*}$ trained the full dataset $S$, thus $\xi_{z^{\prime},\lambda}$ can be calculated without retraining. Moreover, both the LHS and the RHS ($\xi_{z^{\prime},\lambda}$) of Eq. 8 are well normalized.
@@
If $r_{z^{\prime},\lambda}>0$, then LHS and RHS of Eq. 8 both lie in the interval $[0,1]$.
@@
In principle, we need both the LHS and RHS of Eq. 8 to evaluate the sufficient condition in Theorem 4.3. In practice, however, we find that using $\bar{\xi}_{T,\lambda}:=\frac{1}{|T|}\sum_{z^{\prime}\in T}\xi_{z^{\prime},\lambda}=0.5$, where $T$ is a validation dataset, as the criterion [^8] In fact, setting the threshold as 0.4 or 0.6 is similarly effective. The $\xi_{z^{\prime},\lambda}$ as a function of $\log\lambda$ looks like a logistic function and it varies sharply when its value is around 0.5. See Section C.1 for visualizations. to select the regularization term $\lambda$ is uniformly effective across all the experimental settings. This yields a practical algorithm (Algorithm 1) for selecting the hyperparameter $\lambda$ without requiring model retraining.
@@
7:$\hat{\lambda}\leftarrow\operatorname*{arg\,min}_{\lambda\in\mathcal{C}}|\bar{\xi}_{T,\lambda}-0.5|$;
@@ 4.3 Experiments: Validating the Proposed Surrogate-Indicator-Based Selection Algorithm
Dataset, model, and projection dimension: We employ 6 experiment settings, covering different datasets, model scales, and projection dimensions. First two settings apply Logistic Regression models (LR) on MNIST dataset [24], where one utilizes random projection with dimension $512$ and the other uses no projection. They are followed by two settings that apply MLP, with random projection $512$ and $4096$ respectively. Finally, we have two relatively large models, including ResNet-9 on CIFAR-2 dataset [23, 13], and MusicTransformer (MT) on MAESTRO dataset [2, 12]. We fix the random projection dimension to be $4096$ in the last two settings. Attributor: We experiment with the IFFIM attributor defined in Eq. 2 (or Eq. 4 with projection), and the TRAK attributor [28]. Baseline selection method: We also compare the proposed algorithm against a baseline method that selects $\lambda$ based on the spectrum of $F_{S}$ since the main motivation of $\lambda$ is to deal with the singularity of $F_{S}$. Specifically, the baseline method set $\lambda$ to be the eigenvalue of $F_{S}$ at a fixed quantile among $\{10\%,30\%,50\%,70\%,90\%\}$.
Figure 3: The plot of LDS versus $\lambda$ on the TRAK attributor. Please see Figure 2 for the plotting setup.
@@ Results.
We demonstrate the experimental results on the IFFIM and TRAK respectively in Figure 2 and Figure 3. In each subplot, the blue curve represents the change in LDS with respect to $\lambda$. We mark the $\lambda$ chosen by the proposed algorithm by the red solid vertical line, and the baselines in the gray dotted vertical lines. Overall, we observe that the $\lambda$ chosen by the proposed algorithm generally leads to good LDS value that is significantly better than $\lambda=0$, and in many case it is close to the optimal value. Moreover, while the proposed algorithm is based on an analysis of the IFFIM attributor, it generalizes well to the TRAK attributor under various experiment settings with different projection dimensions. In contrast, there is no single fixed quantile for spectrum-based baseline method that consistently performs well, indicating that this naive heuristic is insufficient for selecting $\lambda$.
@@ 5 Discussion and Conclusion
Through extensive experiments across diverse methods, datasets, and models, we demonstrate that many data attribution methods are indeed highly sensitive to hyperparameter configurations, with optimal choices varying significantly across settings. In particular, we show that some hyperparameters—such as regularization and projection dimension—interact in nontrivial ways, amplifying sensitivity and necessitating more careful tuning strategies.
@@ A.1 Hyperparameters Definitions
  • normalization (TracIn) [29]: We may normalize the per-example gradients in TracIn by dividing their L2-norm. In other words, normalization=True means that $\tau_{\mathrm{TracIn}}(z^{\prime},z_{i}):=\sum_{t}\eta_{t}\frac{\nabla_{\theta}f(z^{\prime},\theta_{t})^{\top}}{\|\nabla_{\theta}f(z^{\prime},\theta_{t})^{\top}\|}\frac{\nabla_{\theta}L(z_{i},\theta_{t})}{\|\nabla_{\theta}L(z_{i},\theta_{t})\|}$.
@@ A.3 Computational Resources and Dataset Licenses
Appendix B Omitted Details of the Theoretical Analysis in Section 4
@@ B.1 Relationship between TRAK and IFFIM
$\tau_{\mathrm{TRAK}}(z^{\prime},z_{i}):=\nabla_{\theta}f(z^{\prime},\theta_{S}^{*})^{\top}(\Phi^{\top}R\Phi)^{-1}\nabla_{\theta}f(z_{i},\theta_{S}^{*})\cdot(1-p_{i}),$
where $\Phi\in\mathbb{R}^{n\times p}$ has its $i$th row being $\nabla_{\theta}f(z_{i},\theta_{S}^{*})^{\top}$, $R:=\operatorname{diag}\{p_{i}(1-p_{i})\}_{i=1}^{n}$, and $p_{i}:=p(z_{i},\theta_{S}^{*})$.
Compared to the IFFIM attributor in Eq. 2, there are two main differences. First, the right most gradient term in TRAK is $\nabla_{\theta}f(z_{i},\theta_{S}^{*})(1-p_{i})$ while its counterpart in IFFIM is $-\nabla_{\theta}L(z_{i},\theta_{S}^{*})$. Second, IFFIM has the empiricla FIM $F_{S}$ in the middle while TRAK has $\Phi^{\top}R\Phi$.
@@
We first show that $\nabla_{\theta}f(z_{i},\theta_{S}^{*})(1-p_{i})$ is equivalent to $-\nabla_{\theta}L(z_{i},\theta_{S}^{*})$ as
$-\nabla_{\theta}L(z_{i},\theta_{S}^{*})=-\left.\frac{\partial\ln(1+e^{-f})}{\partial f}\right|_{f=f(z_{i},\theta_{S}^{*})}\nabla_{\theta}f(z_{i},\theta_{S}^{*})=(1-p(z_{i},\theta_{S}^{*}))\nabla_{\theta}f(z_{i},\theta_{S}^{*}).$
@@
B.2 Proof of Theorem 4.3
We prove Theorem 4.3 before Lemma 4.2 for better readability. We first establish several intermediate results for Eq. 8.
@@
where $D_{a}$ is defined in Remark 4.4.
@@
Lemma B.1 facilitates analyzing the variance of $\sum_{z\in A}\tau_{\mathrm{IFFIM},\lambda}(z^{\prime},z)$ in Eq. 6, thereby enabling verification of the following equivalent condition of $\dot{c_{p}}(\lambda;z^{\prime})>0$.
@@
We first simplify $\sum_{z\in A}\tau_{\mathrm{IFFIM},\lambda}(z^{\prime},z)$ in Eq. 6:
@@
where we apply the identity Lemma B.1.
@@
By applying the equality $k$ times on $(X^{\top}X+\lambda I_{p})^{-k}X^{\top}$ to “push $X^{\top}$ through" the inverted matrix, we complete the proof. ∎
Proof of Theorem 4.3.
@@
Without loss of generality, we only consider the case where $-\nabla_{\theta}f(z^{\prime},\theta_{S}^{*})^{\top}(F_{S}+\lambda I_{p})^{-2}g_{z^{\prime}}>0$, because otherwise Eq. 10 holds automatically as its left hand side would be positive while right hand side would be non-positive, which guarantees $\dot{c_{p}}(\lambda;z^{\prime})>0$ by Proposition B.2.
We proceed by rewriting Eq. 10 with Lemma B.3. By setting $X=J/\sqrt{n}$, we have
@@
$-\nabla_{\theta}f(z^{\prime},\theta_{S}^{*})^{\top}(F_{S}+\lambda I_{p})^{-2}g_{z^{\prime}}=-\frac{1}{n}\nabla_{\theta}f(z^{\prime},\theta_{S}^{*})^{\top}J^{\top}(\frac{1}{n}JJ^{\top}+\lambda I_{n})^{-2}\alpha_{z^{\prime}},$
@@
by applying the generalized Cauchy-Schwarz inequality on $-(\frac{1}{n}JJ^{\top}+\lambda I_{n})^{-1}J\nabla_{\theta}f(z^{\prime},\theta_{S}^{*})$ and $\alpha_{z^{\prime}}$ with inner product defined by positive definite matrix $\frac{1}{n}(\frac{1}{n}JJ^{\top}+\lambda I_{n})^{-1}$. Now, we finish the proof by observing that Eq. 10 is derived directly by multiplying Eq. 8 with Eq. 11 and rearranging terms. ∎
B.3 Assumptions and Proof of Lemma 4.2
We base the discussion of Lemma 4.2 on Eq. 8. For simplicity, we first eliminate $\sqrt{t_{1,z^{\prime},\lambda}}$ on both sides to state the following sufficient condition for $\dot{c_{p}}(\lambda;z^{\prime})>0$:
@@
Now, for $i=1,2,\dots,\max\{n,p\}$, we let $\mu_{i}$ denote the $i$th largest eigenvalue of $F_{S}$ if $i\leq p$, and 00 otherwise. We have the following result for $\mu_{i}$:
@@
$C_{1}\mathbb{E}_{z\sim S}[\widetilde{\nabla_{\theta}L}_{i}(z,\theta_{S}^{*})^{2}]\leq\widetilde{\nabla_{\theta}L}_{i}(z^{\prime},\theta_{S}^{*})^{2}\leq C_{2}\mathbb{E}_{z\sim S}[\widetilde{\nabla_{\theta}L}_{i}(z,\theta_{S}^{*})^{2}],$
@@
B.7 basically assumes that the distribution given by $S$ represents the test distribution well regarding the relative size of eigen-components. Then, we are able to derive an upper bound for the RHS (right hand side) of Eq. 12.
@@
Under B.7, for all $\lambda>0$, we have
$\mathrm{RHS}<\frac{C_{2}}{(1-p(z^{\prime},\theta_{S}^{*}))\sqrt{C_{1}}}\frac{(\mu_{\mathrm{min}}+\lambda)^{3/2}}{\mu_{\mathrm{min}}},$ (13)
@@
so the identity holds because $\Sigma_{ii}=\sqrt{n\cdot\mu_{i}}$ for $i\in[\min\{n,p\}]$ (see Lemma B.4) and $\mu_{i}=0$ for $i>p$.
@@
$\leq\frac{1}{(1-p(z^{\prime},\theta_{S}^{*}))^{2}}C_{2}.$
@@
We move to analyze the LHS (left hand side) of Eq. 12. Before that, we need two additional technical assumptions.
@@
$\tilde{g}_{z^{\prime},i}^{2}\leq C_{3}\mu_{i}^{1+\varepsilon_{g}}.$

### v2

@@ 4.2 Analysis: A Sufficient Condition for Monotonicity in $\lambda$
Intuition behind Theorem˜4.3.
Let’s first focus on RHS of Eq.˜8. By observing the definition of $t_{k,z^{\prime},\lambda}$ for $k=1,2,3$, RHS can be viewed as the cosine “angle” between $v_{z^{\prime}}$ and $(F_{S}+\lambda I_{p})^{-1}v_{z^{\prime}}$. Here, $(F_{S}+\lambda I_{p})^{-1}v_{z^{\prime}}$ is related to the $\theta$ update direction in IFFIM, as in Eq.˜2. When RHS is small, this “angle” is overly large, suggesting a significant misalignment between the $\theta$ update direction and the gradient descent direction direction. In this case, increasing $\lambda$ helps intuitively because it conditions the inverted matrix $F_{S}$ better, stabilizing the update direction. When the “angle” is too small, $F_{S}$ is overshadowed by the regularization term $\lambda I_{p}$ (as $v_{z^{\prime}}$ is roughly parallel to $(F_{S}+\lambda I_{p})^{-1}v_{z^{\prime}}$), meaning that the attribution method may become dominated by regularization and lose sensitivity to the underlying structure of data. Similarly, LHS of Eq.˜8 measures the “angle” between $v_{z^{\prime}}$ and the ground-truth direction that involves $g_{z^{\prime}}$. Hence Theorem˜4.3 is in essence stating that when the “angle” between $v_{z^{\prime}}$ and $(F_{S}+\lambda I_{p})^{-1}v_{z^{\prime}}$ is of comparable level as the “angle” between $v_{z^{\prime}}$ and the ground-truth direction, Population Pearson LDS could reach its maximum.
Remark 4.4 (Proof Sketch for Theorem˜4.3).
The detailed proof can be found in Section˜B.3. We provide a proof sketch here. We start by expanding $\dot{c_{p}}(\lambda;z^{\prime})$ with the definition of $\tau_{\mathrm{IFFIM},\lambda}$. Then, We draw connections between $\mathrm{Var}_{A\sim D_{a}}[\nabla_{\theta}R_{A}(\theta_{S}^{*})]$, a crucial term occurred in the expression of $\dot{c_{p}}(\lambda;z^{\prime})$, and $F_{S}$. Further, we state a condition (Eq.˜10) equivalent to $\dot{c_{p}}(\lambda;z^{\prime})>0$. Finally, we derive Eq.˜8 from Eq.˜10 with Lemma˜B.3 and the generalized Cauchy-Schwarz inequality.
@@
A result similar to Theorem˜4.3 can also be derived for IFFIM attributor with gradient projection (see $\tau_{\mathrm{IFFIM},\lambda,P}$ in Eq.˜4). In this case, given the projection matrix $P$, $t_{k,z^{\prime},\lambda}$ should be replaced with $v_{z^{\prime}}^{\top}P(P^{\top}F_{S}P+\lambda I_{\tilde{p}})^{-k}P^{\top}F_{S}PP^{\top}v_{z^{\prime}}$ while $r_{z^{\prime},\lambda}$ should be replaced with $v_{z^{\prime}}^{\top}P(P^{\top}F_{S}P+\lambda I_{\tilde{p}})^{-1}P^{\top}g_{z^{\prime}}$. Additionally, $JJ^{\top}+n\lambda I_{n}$ in the $o_{z^{\prime},\lambda}$ should be replaced with $JPP^{\top}J^{\top}+n\lambda I_{n}$. Please see more details in Section˜B.5.
@@ 4.3 Algorithm: The Surrogate-Indicator-Based Practical Selection
The sufficient condition in Theorem˜4.3 enables a practical selection algorithm for $\lambda$ based on the following surrogate indicator $\xi_{z^{\prime},\lambda}$.
@@
$\xi_{z^{\prime},\lambda}\mathrel{\mathop{\ordinarycolon}}=\frac{t_{2,z^{\prime},\lambda}}{\sqrt{t_{3,z^{\prime},\lambda}\cdot t_{1,z^{\prime},\lambda}}},$ (9)
which is the RHS of Eq.˜8.
Note that the quantities $t_{k,z^{\prime},\lambda}$ for $k=1,2,3$ only depends on model parameters $\theta_{S}^{*}$ trained on the full dataset $S$, thus $\xi_{z^{\prime},\lambda}$ can be calculated without retraining. Moreover, both the LHS and the RHS ($\xi_{z^{\prime},\lambda}$) of Eq.˜8 are well normalized.
@@
If $r_{z^{\prime},\lambda}>0$, then LHS and RHS of Eq.˜8 both lie in the interval $[0,1]$.
@@
In principle, we need both the LHS and RHS of Eq.˜8 to evaluate the sufficient condition in Theorem˜4.3. In practice, however, we find that using $\bar{\xi}_{T,\lambda}\mathrel{\mathop{\ordinarycolon}}=\frac{1}{|T|}\sum_{z^{\prime}\in T}\xi_{z^{\prime},\lambda}=0.5$, where $T$ is a validation dataset, as the criterion [^8] In fact, setting the threshold as 0.4 or 0.6 is similarly effective. The $\xi_{z^{\prime},\lambda}$ as a function of $\log\lambda$ looks like a logistic function and it varies sharply when its value is around 0.5. See Section C.1 for visualizations. to select the regularization term $\lambda$ is uniformly effective across all the experimental settings. This yields a practical algorithm (Algorithm˜1) for selecting the hyperparameter $\lambda$ without requiring model retraining.
@@
[…]
@@ 4.4 Experiments: Validating the Proposed Surrogate-Indicator-Based Selection Algorithm
Dataset, model, and projection dimension: We employ 6 experiment settings, covering different datasets, model scales, and projection dimensions. First two settings apply Logistic Regression models (LR) on MNIST dataset [25], where one utilizes random projection with dimension $512$ and the other uses no projection. They are followed by two settings that apply MLP, with random projection $512$ and $4096$ respectively. Finally, we have two relatively large models, including ResNet-9 [14] on CIFAR-2 dataset [24], and MusicTransformer (MT) [2] on MAESTRO dataset [13]. We fix the random projection dimension to be $4096$ in the last two settings. We also conduct experiments on WikiText2 [28] with GPT2 [31], the results of which can be found in Section˜C.4. Attributor: We experiment with the IFFIM attributor defined in Eq.˜2 (or Eq.˜4 with projection), and the TRAK attributor [29]. Baseline selection method: We also compare the proposed algorithm against a baseline method that selects $\lambda$ based on the spectrum of $F_{S}$ since the main motivation of $\lambda$ is to deal with the singularity of $F_{S}$. Specifically, the baseline method set $\lambda$ to be the eigenvalue of $F_{S}$ at a fixed quantile among $\{10\%,30\%,50\%,70\%,90\%\}$.
Figure 3: The plot of LDS versus $\lambda$ on the TRAK attributor. Please see Figure˜2 for the plotting setup.
@@ Results.
We demonstrate the experimental results on the IFFIM and TRAK respectively in Figure˜2 and Figure˜3. In each subplot, the blue curve represents the change in LDS with respect to $\lambda$. We mark the $\lambda$ chosen by the proposed algorithm by the red solid vertical line, and the baselines in the gray dotted vertical lines. Overall, we observe that the $\lambda$ chosen by the proposed algorithm generally leads to good LDS value that is significantly better than $\lambda=0$, and in many case it is close to the optimal value. Moreover, while the proposed algorithm is based on an analysis of the IFFIM attributor, it generalizes well to the TRAK attributor under various experiment settings with different projection dimensions. In contrast, there is no single fixed quantile for spectrum-based baseline method that consistently performs well, indicating that this naive heuristic is insufficient for selecting $\lambda$.
Downstream tasks performance.
We also include an evaluation of our proposed algorithm on downstream settings, and the detailed results are presented in Section˜C.5. We focus on the task of data selection commonly used in data attribution literature [11], where we measure the performance drop of models when the most positively influential training data are removed. We note here that a larger drop signifies stronger attribution quality. In general, we observe that our proposed algorithm achieves visible model performance drops, demonstrating its effectiveness in downstream tasks.
@@ 5 Discussion and Conclusion
Through extensive experiments across diverse methods, datasets, and models, we demonstrate that many data attribution methods are indeed highly sensitive to hyperparameter configurations, with optimal choices varying significantly across settings. For example, we show that some hyperparameters—such as regularization and projection dimension—interact in nontrivial ways, amplifying sensitivity and necessitating more careful tuning strategies. Moreover, implicit factors (such as training epoch) matter; comparing attributions across checkpoints can be informative.
@@ A.1 Hyperparameters Definitions
  • normalization (TracIn) [30]: We may normalize the per-example gradients in TracIn by dividing their L2-norm. In other words, normalization=True means that $\tau_{\mathrm{TracIn}}(z^{\prime},z_{i})\mathrel{\mathop{\ordinarycolon}}=\sum_{t}\eta_{t}\frac{\nabla_{\theta}f(z^{\prime},\theta_{t})^{\top}}{\|\nabla_{\theta}f(z^{\prime},\theta_{t})^{\top}\|}\frac{\nabla_{\theta}L(z_{i},\theta_{t})}{\|\nabla_{\theta}L(z_{i},\theta_{t})\|}$.
@@ A.3 Computational Resources and Dataset Licenses
Appendix B Omitted Details of the Theoretical Analysis in Section˜4
@@ B.1 Relationship between TRAK and IFFIM
$\tau_{\mathrm{TRAK}}(z^{\prime},z_{i})\mathrel{\mathop{\ordinarycolon}}=\nabla_{\theta}f(z^{\prime},\theta_{S}^{*})^{\top}(\Phi^{\top}R\Phi)^{-1}\nabla_{\theta}f(z_{i},\theta_{S}^{*})\cdot(1-p_{i}),$
where $\Phi\in\mathbb{R}^{n\times p}$ has its $i$th row being $\nabla_{\theta}f(z_{i},\theta_{S}^{*})^{\top}$, $R\mathrel{\mathop{\ordinarycolon}}=\operatorname{diag}\{p_{i}(1-p_{i})\}_{i=1}^{n}$, and $p_{i}\mathrel{\mathop{\ordinarycolon}}=p(z_{i},\theta_{S}^{*})$.
Compared to the IFFIM attributor in Eq.˜2, there are two main differences. First, the right most gradient term in TRAK is $\nabla_{\theta}f(z_{i},\theta_{S}^{*})(1-p_{i})$ while its counterpart in IFFIM is $\nabla_{\theta}L(z_{i},\theta_{S}^{*})$. Second, IFFIM has the empiricla FIM $F_{S}$ in the middle while TRAK has $\Phi^{\top}R\Phi$.
@@
We first show that $\nabla_{\theta}f(z_{i},\theta_{S}^{*})(1-p_{i})$ is equivalent to $\nabla_{\theta}L(z_{i},\theta_{S}^{*})$ as
$\nabla_{\theta}L(z_{i},\theta_{S}^{*})=-\left.\frac{\partial\ln(1+e^{-f})}{\partial f}\right|_{f=f(z_{i},\theta_{S}^{*})}\nabla_{\theta}f(z_{i},\theta_{S}^{*})=(1-p(z_{i},\theta_{S}^{*}))\nabla_{\theta}f(z_{i},\theta_{S}^{*}).$
@@
B.2 LOO Weighted Loss Gradients
As suggested by Park et al. [29], we assume decreasing $a$ by $1$ does not shift the distribution of $f(z^{\prime},\theta_{A}^{*})$, i.e.,
$\mathbb{E}_{A\sim D_{a}}[f(z^{\prime},\theta_{A}^{*})|z_{i}\notin A]=\mathbb{E}_{A\sim D_{a-1}}[f(z^{\prime},\theta_{A}^{*})|z_{i}\notin A].$
Then, by rewriting the RHS, we have
$\mathbb{E}_{A\sim D_{a}}[f(z^{\prime},\theta_{A}^{*})|z_{i}\notin A]=\mathbb{E}_{A\sim D_{a}}[f(z^{\prime},\theta_{A\backslash\{z_{i}\}}^{*})|z_{i}\in A],$
and hence
$\alpha_{z^{\prime},i}$ $=\mathbb{E}_{A\sim D_{a}}[f(z^{\prime},\theta_{A}^{*})|z_{i}\in A]-\mathbb{E}_{A\sim D_{a}}[f(z^{\prime},\theta_{A}^{*})]\vphantom$$\frac${A}{A}}
$=\mathbb{E}_{A\sim D_{a}}[f(z^{\prime},\theta_{A}^{*})|z_{i}\in A]$
$\quad\quad\quad-\Pr_{A\sim D_{a}}[z_{i}\in A]\mathbb{E}_{A\sim D_{a}}[f(z^{\prime},\theta_{A}^{*})|z_{i}\in A]-\Pr_{A\sim D_{a}}[z_{i}\notin A]\mathbb{E}_{A\sim D_{a}}[f(z^{\prime},\theta_{A}^{*})|z_{i}\notin A]$
$=\Pr_{A\sim D_{a}}[z_{i}\notin A](\mathbb{E}_{A\sim D_{a}}[f(z^{\prime},\theta_{A}^{*})|z_{i}\in A]-\mathbb{E}_{A\sim D_{a}}[f(z^{\prime},\theta_{A}^{*})|z_{i}\notin A])$
$=(1-\frac{a}{n})\mathbb{E}_{A\sim D_{a}}[f(z^{\prime},\theta_{A}^{*})-f(z^{\prime},\theta_{A\backslash\{z_{i}\}}^{*})|z_{i}\in A],$
which means that the weights $\alpha_{z^{\prime},i}$ in $g_{z^{\prime}}$ can be interpreted as the LOO influence of $z_{i}$ on $z^{\prime}$.
B.3 Proof of Theorem 4.3
We prove Theorem˜4.3 before Lemma˜4.2 for better readability. We first establish several intermediate results for Eq.˜8.
@@ B.2 LOO Weighted Loss Gradients
where $D_{a}$ is defined in Section˜4.2.
@@
Lemma˜B.1 facilitates analyzing the variance of $\sum_{z\in A}\tau_{\mathrm{IFFIM},\lambda}(z^{\prime},z)$ in Eq.˜6, thereby enabling verification of the following equivalent condition of $\dot{c_{p}}(\lambda;z^{\prime})>0$.
@@
We first simplify $\sum_{z\in A}\tau_{\mathrm{IFFIM},\lambda}(z^{\prime},z)$ in Eq.˜6:
@@
where we apply the identity Lemma˜B.1.
@@
By applying the equality $k$ times on $(X^{\top}X+\lambda I_{p})^{-k}X^{\top}$ to “push $X^{\top}$ through” the inverted matrix, we complete the proof. ∎
Proof of Theorem˜4.3.
@@
Without loss of generality, we only consider the case where $\nabla_{\theta}f(z^{\prime},\theta_{S}^{*})^{\top}(F_{S}+\lambda I_{p})^{-2}g_{z^{\prime}}>0$, because otherwise Eq.˜10 holds automatically as its left hand side would be positive while right hand side would be non-positive, which guarantees $\dot{c_{p}}(\lambda;z^{\prime})>0$ by Proposition˜B.2.
We proceed by rewriting Eq.˜10 with Lemma˜B.3. By setting $X=J/\sqrt{n}$, we have
@@
$\nabla_{\theta}f(z^{\prime},\theta_{S}^{*})^{\top}(F_{S}+\lambda I_{p})^{-2}g_{z^{\prime}}=-\frac{1}{n}\nabla_{\theta}f(z^{\prime},\theta_{S}^{*})^{\top}J^{\top}(\frac{1}{n}JJ^{\top}+\lambda I_{n})^{-2}\alpha_{z^{\prime}},$
@@
by applying the generalized Cauchy-Schwarz inequality on $(\frac{1}{n}JJ^{\top}+\lambda I_{n})^{-1}J\nabla_{\theta}f(z^{\prime},\theta_{S}^{*})$ and $\alpha_{z^{\prime}}$ with inner product defined by positive definite matrix $\frac{1}{n}(\frac{1}{n}JJ^{\top}+\lambda I_{n})^{-1}$. Now, we finish the proof by observing that Eq.˜10 is derived directly by multiplying Eq.˜8 with Eq.˜11 and rearranging terms. ∎
B.4 Assumptions and Proof of Lemma 4.2
We base the discussion of Lemma˜4.2 on Eq.˜8. For simplicity, we first eliminate $\sqrt{t_{1,z^{\prime},\lambda}}$ on both sides to state the following sufficient condition for $\dot{c_{p}}(\lambda;z^{\prime})>0$:
@@
Now, for $i=1,2,\dots,\max\{n,p\}$, we let $\mu_{i}$ denote the $i$th largest eigenvalue of $F_{S}$ if $i\leq p$, and 0 otherwise. We have the following result for $\mu_{i}$:
@@
$C_{1}\mathbb{E}_{z\sim S}[\widetilde{\nabla_{\theta}L}_{i}(z,\theta_{S}^{*})^{2}]\leq\widetilde{\nabla_{\theta}L}_{i}(z^{\prime},\theta_{S}^{*})^{2}\leq C_{2}\mathbb{E}_{z\sim S}[\widetilde{\nabla_{\theta}L}_{i}(z,\theta_{S}^{*})^{2}],$ (13)
@@
Assumption˜B.7 basically assumes that the distribution given by $S$ represents the test distribution well regarding the relative size of eigen-components. Then, we are able to derive an upper bound for the RHS (right hand side) of Eq.˜12.
@@
Under Assumption˜B.7, for all $\lambda>0$, we have
$\mathrm{RHS}<\frac{nC_{2}}{(1-p(z^{\prime},\theta_{S}^{*}))\sqrt{C_{1}}}\frac{(\mu_{\mathrm{min}}+\lambda)^{3/2}}{\mu_{\mathrm{min}}},$ (14)
@@
so the identity holds because $\Sigma_{ii}=\sqrt{n\cdot\mu_{i}}$ for $i\in[\min\{n,p\}]$ (see Lemma˜B.4) and $\mu_{i}=0$ for $i>p$.
@@
$\leq\frac{n}{(1-p(z^{\prime},\theta_{S}^{*}))^{2}}C_{2}.$
@@
We move to analyze the LHS (left hand side) of Eq.˜12. Before that, we need two additional technical assumptions.
@@
$\tilde{g}_{z^{\prime},i}^{2}\leq C_{3}\mu_{i}^{1+\varepsilon_{g}}.$ (15)